One right triangle has sides of length 3 cm, 4 cm and 5 cm. Another right triangle
has two sides of length 3 cm and 4 cm, but it is not congruent to the first triangle. What is the length of the third side of this triangle? Express your answer in
simplest radical form

Answer by richard1234(7193)

The hypotenuse is 4 cm (it cannot be 3 cm, and if they were both legs, the hypotenuse would be 5 cm).

Therefore by Pythagorean theorem, a^2 + (3 cm)^2 = (4 cm)^2, so a = sqrt(7) cm.